OpenStudy (anonymous):

Solve the following system by elimination. 3x - 2y = 18 5x + 2y = 14 A. (0, 7) B. (4, -3) C. (6, -6) D. (2, 2)

OpenStudy (anonymous):

\[8x = 32\]\[x=4\]

OpenStudy (anonymous):

There is only one pair with \(x=4\) so this is enough to choose the correct answer.

OpenStudy (nurali):

3x - 2y = 18 Eq. 1 + 5x + 2y = 14 Eq. 2 ------------------------- 8x = 32 (-2y from eq. 1 and 2y from eq. 2 are add. inverses) x = 4 (divide both sides of equation by 8) To get value for y, substitute x = 4 for one of the equations. I 3(4) - 2y = 18 12 - 2y = 18 -2y = 6 (subtract 12 from both sides of equation) y = -3 (divide by -2 on both sides of equation) x= 4, y= -3